# The dark matter content of Milky Way dwarf spheroidal galaxies: Draco, Sextans and Ursa Minor

### v1

@@ 4.4 Astrophysical factors
In this subsection, we provide our measurements of the astrophysical factors of Draco, Sextans and UMi, for the purposes of future dark matter indirect detections (e.g. Baltz et al., 1999; Bergström & Hooper, 2006; Geringer-Sameth et al., 2015; Hayashi et al., 2016). The astrophysical factor is determined by the dark matter distribution within the galaxy. For dark matter annihilation or decay, the astrophysical factor is called the J or D factor, respectively. The J and D factors can be written as
$J$ $=\iint\rho_{\mathrm{DM}}^{2}(l,\Omega)\,\mathrm{d}l\mathrm{d}\Omega$ (20)
$,D$ $=\iint\rho_{\mathrm{DM}}(l,\Omega)\,\mathrm{d}l\mathrm{d}\Omega,$ (21)
@@
Object $\log_{10}J(0.5^{\circ})$ $\log_{10}D(0.5^{\circ})$
$\left[\mathrm{GeV^{2}\,cm^{-5}}\right]$ $\left[\mathrm{GeV\,cm^{-2}}\right]$
Draco $18.74^{+0.12}_{-0.09}$ $18.07^{+0.10}_{-0.10}$
Sextans $17.98^{+0.12}_{-0.12}$ $18.03^{+0.07}_{-0.06}$
UMi $18.27^{+0.07}_{-0.07}$ $18.20^{+0.04}_{-0.05}$
Table 6: Values of J and D factors for three dSphs within $0.5^{\circ}$.
The J and D factors of the three dSphs are provided in Table 6, with integrals in Equation 21 over a fixed solid angle of $\Delta\Omega=0.5^{\circ}$. In Figures 11 and 12, J and D factors calculated from Hayashi et al. (2020) and this work are compared. The differences are mainly due to different dark matter halo models adopted and different truncation radii of dark matter halos. Hayashi et al. (2020) assumed axisymmetric dark matter halos with generalized Hernquist profiles (Hernquist, 1990; Zhao, 1996), while in our analysis we consider spherical dark matter halos with generalized NFW profiles, and the outer density slope has been fixed to 4. On the other hand, both Hayashi et al. (2020) and us truncate the dark matter halo at the projected radius of the outermost observed member star for each dSphs. However, stars beyond $2.5$ half-light radius from the center of each dSphs are not used in our analysis (see Section 2.4), and thus the truncation radius for the dark matter halo in our analysis, i.e., $2.5$ half-light radius, is smaller than that of Hayashi et al. (2020). In Figures 11 and 12, the most prominent difference between Hayashi et al. (2020) and this work is the D factor for Draco. This is because their truncation radius for Draco is much larger than ours and the dark matter distribution at larger radii contributes more to the D factor than the J factor. Besides, since we truncate dark matter halos at smaller radii where dark matter density profiles are better constrained, uncertainties of our J and D factors are also reduced. In our results, Draco turns out to be the most promising target to search for signals of dark matter annihilation and decay.
Figure 11: Measured J factors for three dSphs. Green and magenta symbols denote those of Hayashi et al. (2020) and this work.
Figure 12: Measured D factors for three dSphs. Green and magenta symbols denote those of Hayashi et al. (2020) and this work.
5 Discussions
In Read et al. (2019), the dark matter density at 150 pc, $\rho_{\mathrm{DM}}(150\mathrm{pc})$, is used as a proxy of the inner density slope $\gamma$. We compare our results of $\rho_{\mathrm{DM}}(150\mathrm{pc})$ with those reported by Read et al. (2019) and Hayashi et al. (2020) in Figure 13. For the three dSphs used in our analysis, we mark our best recovered $\rho_{\mathrm{DM}}(150\mathrm{pc})$ from the chemodynamical model above by the magenta filled circles, and we do not repeatedly show those for the single-population model as they are very similar (see Figure 8). Previous measurements by Read et al. (2019) and Hayashi et al. (2020) are plotted with black stars and green diamonds.
@@ 5 Discussions
For Draco, our measured $\rho_{\mathrm{DM}}(150\mathrm{pc})$ agrees well with both Read et al. (2019) and Hayashi et al. (2020). For Sextans, three measurements are different from each other, though the measurements of Hayashi et al. (2020) and Read et al. (2019) still agree within the errors. This discrepancy is very likely due to the fact that both Hayashi et al. (2020) and us adopt axisymmetric Jeans analysis, while Read et al. (2019) used spherical Jeans models. Thus our measurement and the measurement by Hayashi et al. (2020) for Sextans are both lower than that of Read et al. (2019). Besides, the sample of member stars used in Hayashi et al. (2020) is different from that in this work. We have carefully repeated our calculation by using the same sample as in Hayashi et al. (2020), and we can reproduce the result of Hayashi et al. (2020) for Sextans, as shown by the hollow magenta triangle in Figure 13. For UMi, our result is inconsistent with both Read et al. (2019) and Hayashi et al. (2020), and the difference is larger than that for Sextans. We have explicitly checked that this disagreement is mainly due to the large difference in the sample of member stars used for UMi, as the member star sample with LOSV measurements in our analysis is three times larger than those in Hayashi et al. (2020) or Read et al. (2019). We have also repeated our analysis by using the same sample of UMi as in Hayashi et al. (2020), and we again obtain consistent result with both Read et al. (2019) and Hayashi et al. (2020), with the best constraint shown by the hollow magenta triangle in Figure 13. This test robustly confirms that it is mainly the difference in the sample of member stars used in different analysis that leads to the discrepancy between our measurement and those in the two previous studies for UMi.
@@
We have checked that the intrinsic velocity dispersion profiles for the member star sample of Hayashi et al. (2020) and of our analysis above (DESI $+$ Walker et al. (2023)) differ for both Sextans and UMi, and in turn significantly influence the inferred dark matter content. Here the intrinsic velocity dispersions deconvolved from observational errors are calculated following the method [^7] They use a two-parameter Gaussian likelihood function similar to that of Walker et al. (2006) to calculate the systemic velocity and the intrinsic velocity dispersion of a sample of stars: $\log\mathcal{L}=-\frac{1}{2}\left[\sum\log(\sigma_{v_{\text{sys}}}^{2}+\sigma_{v_{i}}^{2})+\sum(v_{i}-v_{\text{sys}})^{2}/(\sigma_{v_{i}}^{2}+\sigma_{v_{\text{sys}}}^{2})\right]$, where $v_{\text{sys}}$ and $\sigma_{v_{\text{sys}}}$ are the systemic velocity and the intrinsic velocity dispersion of the sample, and $v_{i}$ and $\sigma_{v_{i}}$ are the velocity and the velocity uncertainty of each star. in Li et al. (2017), and we have corrected LOSV offsets between the two data samples. We find that the sample in Hayashi et al. (2020) has higher intrinsic velocity dispersions in the central region than our sample, and this is the reason for the lower density $\rho_{\mathrm{DM}}(150\mathrm{pc})$ obtained in our analysis. In fact, the member stars of UMi in Hayashi et al. (2020) are more metal-poor than our sample, which results in higher velocity dispersion in the center. However, given the significantly increased number of DESI observed member stars for UMi, we believe our sample is more complete.
@@
Given the difference in the intrinsic velocity dispersions based on DESI and Walker et al. (2023) member stars for Draco and Sextans, we repeat our modeling using only DESI data for them. We show the dark matter density profiles in Figure 15 of Appendix A. For Draco and Sextans, indeed the best constrained dark matter density profiles differ when we use only DESI observed member stars, and when using the combined sample of DESI $+$ Walker et al. (2023). Nevertheless, the differences are still marginally consistent within the large 1-$\sigma$ errors.
The best recovered $\rho_{\mathrm{DM}}(150\mathrm{pc})$ of three dSphs using only DESI data are shown by the blue filled circles in Figure 13. The intrinsic velocity dispersions of Sextans/Draco, if only using DESI observed stars, are higher/lower in the central region than using the combined data, hence resulting in higher/lower $\rho_{\mathrm{DM}}(150\mathrm{pc})$. For Sextans, since the difference between the blue and magenta circles of Sextans is less significant compared with the errorbars, we think selection effects associated with Sextans are not strong enough to violate the conclusion. For Draco, DESI is missing a lot of metal-poor stars in the central region, which is at least partially because the number of observed member stars in DESI is small (less than 100, see Table 2). On the other hand, Walker et al. (2023) observes more member stars in the center of Draco. Hence the best constrained $\rho_{\mathrm{DM}}(150\mathrm{pc})$ for Draco is more robust with the combined data set of DESI $+$ Walker et al. (2023), i.e., the magenta circle in Figure 13.
Figure 14: The dark matter densities at $150\mathrm{pc}$ ($y$-axis) versus the stellar-to-halo mass ratios ($x$-axis) for dwarf galaxies in a few previous studies. The black filled squares and stars denote results in Read et al. (2019) for dwarf irregular galaxies (dIrrs) and dwarf spheroidal galaxies (dSphs), respectively. The green symbols represent results in Hayashi et al. (2020). The magenta filled circles represent results in our main analysis using the combined member star sample from DESI observation and Walker et al. (2023), based on the best constrained chemodynamical model in this work.
@@
Even with only three dSphs, we find a diversity of $\rho_{\mathrm{DM}}(150\mathrm{pc})$, or the inner density slope, $\gamma$, with $\rho_{\mathrm{DM}}(150\mathrm{pc})$ spanning one order of magnitude for the three dSphs. Draco has the highest density $\rho_{\mathrm{DM}}(150\mathrm{pc})$ with the inner density slope $\gamma=0.71^{+0.34}_{-0.35}$ close to $1$, while dark halos of Sextans and UMi are less dense in inner regions with smaller $\gamma$, $0.26^{+0.22}_{-0.12}$ and $0.33^{+0.20}_{-0.16}$, which are closer to cored models. Moreover, by comparing the measured inner densities for the three dSphs used in our analysis with Read et al. (2019) and Hayashi et al. (2020), and by repeating our analysis using member stars observed in different surveys, we have demonstrated that the currently best constrained inner density slopes are still sensitive to the model assumptions and to sample selection effects.
@@ 6 Conclusions
We find the best recovered model parameters and the underlying dark matter radial density profiles are consistent between single-population and chemodynamical models. The dark matter density profiles are best constrained around the half-number radii of tracer member stars. A diversity is found for the inner density slopes $\gamma$ of host dark matter halos for the three dSphs. The inner density slopes are $0.71^{+0.34}_{-0.35}$, $0.26^{+0.22}_{-0.12}$ and $0.33^{+0.20}_{-0.16}$ for Draco, Sextans and UMi, respectively.
@@
With the chemodynamical model, member stars with spectroscopic observations of each dSph are divided into metal-rich and metal-poor populations, based on their chemical, spatial and dynamical properties. The metal-rich populations are more centrally concentrated, while the metal-poor populations are more extended. Rotations around minor axes are found for both populations. Moreover, the metal-rich populations are dynamically colder, featuring lower velocity dispersion profiles than the metal-poor populations in all three dSphs.
@@
We compare our results with those in two important previous studies, Read et al. (2019) and Hayashi et al. (2020). Our measurement for Draco agrees well with both of Read et al. (2019) and Hayashi et al. (2020). For Sextans, all three measurements are different from each other. This discrepancy is mainly because both Hayashi et al. (2020) and us adopt axisymmetric Jeans analysis, whereas the model of Read et al. (2019) is based on spherical Jeans modeling. Besides, the sample of member stars used by Hayashi et al. (2020) has higher velocity dispersion in the center, resulting in higher dark matter density. Our best recovered $\rho_{\mathrm{DM}}(150)$ of UMi is much lower than both Read et al. (2019) and Hayashi et al. (2020), which is because the UMi member stars in Read et al. (2019) and Hayashi et al. (2020) are more metal-poor and have higher velocity dispersion in the center than ours. Since the member star sample observed by DESI is significantly larger than those in previous observations, especially for UMi, we believe our sample subjects less to selection effects.
@@
In addition to UMi, we also see possible sample selection effects for Draco and Sextans, i.e., their intrinsic velocity dispersions based on the sample of member stars in Walker et al. (2023) are different from those based on only DESI observed member stars. We thus compare the best constraints of using only DESI member stars with the DESI $+$ Walker et al. (2023) data. For Sextans, the best constrained $\rho_{\mathrm{DM}}(150\mathrm{pc})$ by using DESI data is slightly higher than that of using the combined data set, but the difference is less significant compared with errors. For Draco, if only using DESI observed member stars, we find $\rho_{\mathrm{DM}}(150\mathrm{pc})$ becomes significantly lower than that of using DESI $+$ Walker et al. (2023) data, mainly due to the the lacking of DESI observed metal-poor member stars in central regions.
@@
We also estimate astrophysical J and D factors of the three dSphs. Draco has the largest J or D factors, which makes it the most promising target for searches of dark matter annihilation and decay signals.
The data used for all figures in this work can be found at https://zenodo.org/records/15723563.

### v2

@@ 4.4 Astrophysical factors
In this subsection, we provide our measurements of the astrophysical factors of Draco, Sextans and UMi, for the purposes of dark matter indirect detections, such as Fermi-LAT (Atwood et al., 2009), DAMPE (Chang et al., 2017), HAWC (Abeysekara et al., 2017), CTA (Actis et al., 2011) and LHAASO (Cao et al., 2019). The astrophysical factor is determined by the dark matter distribution within the galaxy (e.g. Baltz et al., 1999; Bergström & Hooper, 2006; Geringer-Sameth et al., 2015; Hayashi et al., 2016; Strigari, 2018). For dark matter annihilation or decay, the astrophysical factor is called the J or D factor, respectively. The J and D factors can be written as
$J$ $=\iint\rho_{\mathrm{DM}}^{2}(l,\Omega)\,\mathrm{d}l\mathrm{d}\Omega,$ (20)
$D$ $=\iint\rho_{\mathrm{DM}}(l,\Omega)\,\mathrm{d}l\mathrm{d}\Omega,$ (21)
@@
We calculate the J and D factors of the three dSphs with integrals covering a fixed solid angle of $\Delta\Omega=0.5^{\circ}$, and compare our results with those from Hayashi et al. (2020) in Figure 11. The differences are mainly due to different dark matter halo models adopted and different truncation radii. Hayashi et al. (2020) assumed axisymmetric dark matter halos with generalized Hernquist profiles (Hernquist, 1990; Zhao, 1996), whereas we consider spherical dark matter halos with generalized NFW profiles, and the outer density slopes are fixed to 4. Both Hayashi et al. (2020) and us truncate the dark matter halo at the projected radius of the outermost observed member star for each dSphs. However, stars beyond $2.5$ half-light radius from the center of each dSphs are not used in our analysis, and thus the truncation radius in our analysis is smaller than that of Hayashi et al. (2020). In Figure 11, the most prominent difference between Hayashi et al. (2020) and this work is the D factor for Draco. This is because their truncation radius for Draco is much larger than ours and the dark matter distribution at larger radii contributes more to the D factor than the J factor. Besides, since we truncate at smaller radii where dark matter density profiles are better constrained, uncertainties of our J and D factors are smaller.
Figure 11: Measured J (top) and D (bottom) factors for the three dSphs. Green and magenta symbols denote those from Hayashi et al. (2020) and this work.
5 Discussion
In Read et al. (2019), the dark matter density at 150 pc, $\rho_{\mathrm{DM}}(150\mathrm{pc})$, is used as a proxy of the inner density slope $\gamma$. We compare our results of $\rho_{\mathrm{DM}}(150\mathrm{pc})$ with those reported by Read et al. (2019) and Hayashi et al. (2020) [^6] We use the updated values kindly provided by Kohei Hayashi, which are different from the values listed in Table 2 of his work. in Figure 12. For the three dSphs used in our analysis, we mark our best recovered $\rho_{\mathrm{DM}}(150\mathrm{pc})$ from the chemodynamical model above by the magenta filled circles, and we do not repeatedly show those for the single-population model as they are very similar (see Figure 8). Previous measurements by Read et al. (2019) and Hayashi et al. (2020) are plotted with black stars and green diamonds.
@@ 5 Discussion
For Draco, our measured $\rho_{\mathrm{DM}}(150\mathrm{pc})$ agrees well with both Read et al. (2019) and Hayashi et al. (2020). For Sextans, three measurements are different from each other, though the measurements of Hayashi et al. (2020) and Read et al. (2019) still agree within the errors. This discrepancy is very likely due to the fact that both Hayashi et al. (2020) and us adopt axisymmetric Jeans analysis, while Read et al. (2019) used spherical Jeans models. Thus our measurement and the measurement by Hayashi et al. (2020) for Sextans are both lower than that of Read et al. (2019). Besides, the sample of member stars used in Hayashi et al. (2020) is different from that in this work. We have carefully repeated our calculation by using the same sample as in Hayashi et al. (2020), and we can reproduce the result of Hayashi et al. (2020) for Sextans, as shown by the hollow magenta triangle in Figure 12. For UMi, our result is inconsistent with both Read et al. (2019) and Hayashi et al. (2020), and the difference is larger than that for Sextans. We have explicitly checked that this disagreement is mainly due to the large difference in the sample of member stars used for UMi, as the member star sample with LOSV measurements in our analysis is three times larger than those in Hayashi et al. (2020) or Read et al. (2019). We have also repeated our analysis by using the same sample of UMi as in Hayashi et al. (2020), and we again obtain consistent result with both Read et al. (2019) and Hayashi et al. (2020), with the best constraint shown by the hollow magenta triangle in Figure 12. This test robustly confirms that it is mainly the difference in the sample of member stars used in different analyses that leads to the discrepancy between our measurement and those in the two previous studies for UMi.
@@
We have checked that the intrinsic velocity dispersion profiles for the member star sample of Hayashi et al. (2020) and of our analysis above (DESI ++ Walker et al. (2023)) differ for both Sextans and UMi, and in turn significantly influence the inferred dark matter content. Here the intrinsic velocity dispersions deconvolved from observational errors are calculated following the method [^7] They use a two-parameter Gaussian likelihood function similar to that of Walker et al. (2006) to calculate the systemic velocity and the intrinsic velocity dispersion of a sample of stars: $\log\mathcal{L}=-\frac{1}{2}\left[\sum\log(\sigma_{v_{\text{sys}}}^{2}+\sigma_{v_{i}}^{2})+\sum(v_{i}-v_{\text{sys}})^{2}/(\sigma_{v_{i}}^{2}+\sigma_{v_{\text{sys}}}^{2})\right]$, where $v_{\text{sys}}$ and $\sigma_{v_{\text{sys}}}$ are the systemic velocity and the intrinsic velocity dispersion of the sample, and $v_{i}$ and $\sigma_{v_{i}}$ are the velocity and the velocity uncertainty of each star. in Li et al. (2017), and we have corrected LOSV offsets between the two data samples. We find that the sample in Hayashi et al. (2020) has higher intrinsic velocity dispersions in the central region than our sample, and this is the reason for the lower density $\rho_{\mathrm{DM}}(150\mathrm{pc})$ obtained in our analysis. In fact, the member stars of UMi in Hayashi et al. (2020) are more metal-poor than our sample, which results in higher velocity dispersion in the center. However, given the significantly increased number of DESI observed member stars for UMi, we believe our sample is more complete.
@@
Given the difference in the intrinsic velocity dispersions based on DESI and Walker et al. (2023) member stars for Draco and Sextans, we repeat our modeling using only DESI data for them. For Draco and Sextans, the best constrained dark matter density profiles slightly differ when we use only DESI observed member stars, and when using the combined sample of DESI ++ Walker et al. (2023). Nevertheless, the differences are still marginally consistent within the large 1-$\sigma$ errors.
The best recovered $\rho_{\mathrm{DM}}(150\mathrm{pc})$ of three dSphs using only DESI data are shown by the blue filled circles in Figure 12. Using only DESI observed stars yields higher intrinsic velocity dispersions in the central region for Sextans whereas lower ones for Draco than using the combined data, hence resulting in higher and lower $\rho_{\mathrm{DM}}(150\mathrm{pc})$, respectively. For Sextans, since the difference between the blue and magenta circles is less significant compared with the errorbars, we think selection effects associated with Sextans are not strong enough to violate the conclusion. For Draco, DESI is missing a lot of metal-poor stars in the central region, which is at least partially because the number of observed member stars in DESI is small (less than 100, see Table 1). On the other hand, Walker et al. (2023) observes more member stars in the center of Draco. Hence the best constrained $\rho_{\mathrm{DM}}(150\mathrm{pc})$ for Draco is more robust with the combined data set of DESI ++ Walker et al. (2023), i.e., the magenta circle in Figure 12.
Figure 13: The dark matter densities at $150\mathrm{pc}$ ($y$-axis) versus the stellar-to-halo mass ratios ($x$-axis) for dwarf galaxies in a few previous studies. The black filled squares and stars denote results in Read et al. (2019) for dwarf irregular galaxies (dIrrs) and dwarf spheroidal galaxies (dSphs), respectively. The green symbols represent results in Hayashi et al. (2020). The magenta filled circles represent results in our main analysis using the combined member star sample from DESI observation and Walker et al. (2023), based on the best constrained chemodynamical model in this work. No obvious correlation is observed.
@@
Even with only three dSphs, we find a diversity of $\rho_{\mathrm{DM}}(150\mathrm{pc})$, or the inner density slope, $\gamma$, with $\rho_{\mathrm{DM}}(150\mathrm{pc})$ spanning nearly a factor of five among the three dSphs. Draco has the highest density $\rho_{\mathrm{DM}}(150\mathrm{pc})$ with the inner density slope $\gamma=0.71^{+0.34}_{-0.35}$ close to $1$, while dark halos of Sextans and UMi are less dense in inner regions with smaller $\gamma$, $0.26^{+0.22}_{-0.12}$ and $0.33^{+0.20}_{-0.16}$, which are closer to cored models. Moreover, by comparing the measured inner densities with Read et al. (2019) and Hayashi et al. (2020), and by repeating our analysis using member stars observed in different surveys, we have demonstrated that the currently best constrained inner density slopes are still sensitive to the model assumptions and to sample selection effects.
@@ 6 Conclusions
We find the best recovered model parameters and the underlying dark matter radial density profiles are consistent between single-population and chemodynamical models. The dark matter density profiles are best constrained around the half-number radii of tracer member stars. A diversity is found for the inner density slopes $\gamma$ of host dark matter halos. The inner density slopes are $0.71^{+0.34}_{-0.35}$, $0.26^{+0.22}_{-0.12}$ and $0.33^{+0.20}_{-0.16}$ for Draco, Sextans and UMi, respectively.
@@
With the chemodynamical model, member stars with spectroscopic observations of each dSph are divided into metal-rich and metal-poor populations, based on their chemical, spatial and dynamical properties. The metal-rich populations are more centrally concentrated, while the metal-poor populations are more extended. Moreover, the metal-rich populations are dynamically colder, featuring lower velocity dispersion profiles than the metal-poor populations in all three dSphs. We identify large differences between the velocity dispersions of the metal-rich and metal-poor populations, likely indicating deviations from the steady-state model assumption.
We identify rotations in all three dSphs. In particular, the metal-rich and metal-poor populations in Draco show the same sign of rotation around the minor axes but with different amplitudes. The two populations in Sextans show different signs of rotations. For UMi, rotation is only seen in the metal-poor population.
We still observe some correlations between the velocity anisotropy parameter and the potential parameters, indicating the existence of mass-anisotropy degeneracy in our results. The Gaia proper motions are subject to very large uncertainties at the distances of our three dSphs, and thus including or not including Gaia proper motions lead to almost the same constraints. The mass-anisotropy degeneracy can be broken with more precise proper motion measurements in the future, or with higher order Jeans analysis (e.g. Merrifield & Kent, 1990; Łokas, 2002; Richardson & Fairbairn, 2013; Wardana et al., 2025).
@@
We compare our results with those in two important previous studies, Read et al. (2019) and Hayashi et al. (2020). Our measurement for Draco agrees well with both of Read et al. (2019) and Hayashi et al. (2020). For Sextans, all three measurements are different from each other. This discrepancy is mainly because both Hayashi et al. (2020) and this study adopt axisymmetric Jeans analysis, whereas the model of Read et al. (2019) is based on spherical Jeans modeling. Besides, the sample of member stars used by Hayashi et al. (2020) has higher velocity dispersion in the center, resulting in higher dark matter density. Our best recovered $\rho_{\mathrm{DM}}(150)$ of UMi is much lower than both Read et al. (2019) and Hayashi et al. (2020), which is because the UMi member stars in Read et al. (2019) and Hayashi et al. (2020) are more metal-poor and have higher velocity dispersion in the center than ours. Since the member star sample observed by DESI is significantly larger than those in previous observations, especially for UMi, we believe our sample subjects less to selection effects.
@@
In addition to UMi, we also see possible sample selection effects for Draco and Sextans, i.e., their intrinsic velocity dispersions based on the sample of member stars in Walker et al. (2023) are different from those based on only DESI observed member stars. We thus compare the best constraints of using only DESI member stars with the DESI ++ Walker et al. (2023) data. For Sextans, the best constrained $\rho_{\mathrm{DM}}(150\mathrm{pc})$ by using DESI data is slightly higher than that of using the combined data set, but the difference is less significant compared with errors. For Draco, if only using DESI observed member stars, we find $\rho_{\mathrm{DM}}(150\mathrm{pc})$ becomes significantly lower than that of using DESI ++ Walker et al. (2023) data, mainly due to the the lacking of DESI observed metal-poor member stars in central regions.
@@
We also estimate astrophysical J and D factors of the three dSphs. The data used for all figures in this work, including the best-fit inner densities and the J and D factors, can be found at https://zenodo.org/records/16919342.